OpenStudy (anonymous):

Find the midpoint of each segment with the given end points, C (1, -5) and D (-5, 1). a. (-2, -3) b. (-2, -2) c. (3, -2) d. (3, -3)

OpenStudy (anonymous):

b

OpenStudy (angela210793):

XM=(XC+XD)/2 and YM=(YC+YD)/2